OpenStudy (darryqx):

Okay, I am stuck in this geometry question that actually seems pretty simple. Am I using the wrong method?

OpenStudy (darryqx):

http://prntscr.com/arzpjg

OpenStudy (darryqx):

See, I tried to find the value for x, and got 7.2, but apparently, that is incorrect. Am I using the wrong method to solve this?

OpenStudy (anonymous):

i can see why you are stuck it is not clear from that picture which side is largest

OpenStudy (anonymous):

maybe try \[\frac{x}{6}=\frac{24}{20}\]

OpenStudy (anonymous):

i get \[x=\frac{36}{5}\] but that is just my guess

OpenStudy (anonymous):

oh, that is \(7.2\) and you said it was wrong, damn

OpenStudy (darryqx):

Yes, I see. I would assume it's segment BC, but it doesn't really state it. And yes, I did that equation and it came out to the same answer. I'm not sure why it would be incorrect, but it is.

OpenStudy (anonymous):

only other guess i would have is \[\frac{x}{6}=\frac{20}{24}\]

OpenStudy (anonymous):

in that case you get \(x=5\) it has to be one of those two for sure!

OpenStudy (anonymous):

OOH i see !

OpenStudy (anonymous):

it says ABC and DEF in that order, so yes, the second one

OpenStudy (darryqx):

Ah, I see. Thank you so much, I really do appreciate it.

OpenStudy (anonymous):

yw